OpenStudy (anonymous):

The measures of the legs of a triangle are 5 inches, 6 inches, and 10 inches. Which of these explains why the triangle is scalene?

OpenStudy (anonymous):

A scalene triangle is a triangle that has three unequal sides.

OpenStudy (anonymous):

what percents is equivalent to 18/25?

OpenStudy (anonymous):

18/25=0.72=72%

OpenStudy (anonymous):

wich fraction is equivalent to 48/5

OpenStudy (anonymous):

\[\left\{9+\frac{3}{5},\frac{96}{10}\right\} \]